log 3log 9+log 729+3log √3³ find the value

Grade:12

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student
3log3*3=6log3 729=3^6
log729=6log3
log v3³=(3/2)log3=1.5log3
3log 9+log 729+3log v3³=6log3+6log3+4.5log3=16.5log3
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty